OpenStudy (anonymous):

Can someone help me please? find the limit of (x)/(e^(1/x) +1) for when x->0

OpenStudy (anonymous):

\[\lim x->0 for when (\frac{ x }{ e ^{\frac{ 1 }{ x }} +1})\]

OpenStudy (anonymous):

the denominator goes to infinity lickety split the numerator goes to zero

OpenStudy (anonymous):

do i need to do any algebra or is it pure speculation?

OpenStudy (anonymous):

not sure what algebra you need here the fact that \[\lim_{x\to 0}x=0\] is more or less like whose buried in grant's tomb

OpenStudy (anonymous):

as for \(e^{\frac{1}{x}}\) that limit is \(\infty\) if \(x\to 0^+\) as \(\lim_{x\to 0^+}\frac{1}{x}=\infty\)

OpenStudy (anonymous):

if \(x\to 0^-\) then \(\lim_{x\to 0^-}\frac{1}{x}=-\infty\) and so \[\lim_{x\to 0^-}e^{\frac{1}{x}}=0\]

OpenStudy (anonymous):

oh i get it now.... thanks a bunch :D

OpenStudy (anonymous):

yw